OpenStudy (anonymous):

HELP which expression is equivalent to (b+h) ( b+h) ?? A .B^2 + 2BH +H^2 B. b^2 -B^2 C. B^2 -2BH - H^2 D. b^2 + 2bh h^

OpenStudy (anonymous):

A:

OpenStudy (anonymous):

positive :)?

OpenStudy (anonymous):

\[(b+h)^2=(b+h)(b+h)=b{}^{\wedge}2+b h+h b +h{}^{\wedge}2= \]\[b{}^{\wedge}2+b h+b h +h{}^{\wedge}2=b{}^{\wedge}2+2b h+h{}^{\wedge}2 \]

OpenStudy (anonymous):

The correct answer is C.

OpenStudy (anonymous):

How could it be C.? There are no negative signs in the problem statement.

OpenStudy (anonymous):

Oh im sorry i didnt see them. Well as B and H are supossed to be positive number then the correct answer is A.

OpenStudy (anonymous):

I thing you are right.

OpenStudy (anonymous):

It's not like they are positive or negative, A B and H are just generic real numbers, they could be either positive or negative.

OpenStudy (anonymous):

Oops, sorry! There is no number A hehe.